OpenStudy (anonymous):

how to solve this using Cauchy Euler x^2y''+xy'+y=0

OpenStudy (amistre64):

decide what x^m you want to apply to it

OpenStudy (amistre64):

spose y = x^m y' = mx^(m-1) y''= m(m-1)x^(m-2) sub those into the eq to find values for m; or call them r for roots if you want